sadiemae:

Use the information to write an equation with variables on both sides. Let x represent Kim’s age. Bob is 4 years older than twice Kim’s age. Paul is 1 year younger than 3 times Kim's age. Bob and Paul are the same age. Which equation can be used to determine Kim’s age, x, from the information about Bob and Paul? 4x + 2 = 3x + 1 4x + 2 = 3x + (–1) 2x + 4 = 3x + 1 2x + 4 = 3x + (–1)

sadiemae:

I'm stuck

darkknight:

okay, since bob and paul are the same age, lets just assign both to the variable y. and K as the variable for kim So we know that 2k+4 = y and that 3k-1 = y we can now use substitution, and say that 3k-1 = 2k+4 Now the problem uses x instead of k, so replace k with x, and you should get your answer :)